# A Survey on Large Language Model-Based Game Agents

### v3

@@ 6. Multi-LLMGA Framework
In this section, we extend the single agent framework to multi-agent settings. Designing a multi-agent system in games is different from generic multi-agent system formulations because games impose unique constraints: states are rarely fully observable, communication channels are often bandwidth-limited, and in certain scenarios direct memory sharing is disallowed to preserve realistic simulation. To analyze how existing work addresses these challenges, we distinguish two complementary dimensions of design.
At the agent level, we examine how agents exchange information and integrate it into their decision-making. Communication protocols specify what message is transmitted (e.g., observations, beliefs, or intentions) and how it is interpreted by receivers. At the organization level, we study how agents are organized as a collective. Organizational structure encompasses three aspects: the topology of connections that shape decision flow, the allocation of tasks and roles that governs functional division of labor, and the mechanisms for ensuring scalability and robustness as groups expand. Figure 8 presents the structure of this section of different components within the multi-agent system.
Figure 8. Overview of the multi-agent framework.
@@ 6.2. Organizational Structure
Organizational structure defines how agents are arranged and coordinated within a multi-agent organization, including the topology of their connections, the allocation of tasks and roles, and the mechanisms that ensure scalability and stability as the population grows.
@@ 8.4. Multi-LLMGA System
Prior studies have demonstrated the potential of multi-agent systems in large-scale simulations, where agents exhibit emergent behaviors. However, current large-scale multi-agent simulations remain constrained by structural and methodological limitations. Many emergent phenomena, such as role differentiation, norm formation, or collective planning, are closely tied to task initialization and rule design. In practice, agents are often seeded with shared goals, cooperation-oriented prompts, or predefined role templates that guide subsequent division of labor and coordination patterns. Prior studies of multi-agent societies have shown that such structural priors are widespread, from small-scale social environments (park2023generative, ) to hierarchical and large-scale simulations (sagents, ; zhao2024hierarchical, ; sid, ), where coordination often reflects the constraints of task setup rather than fully autonomous self-organization. Moreover, the lack of open and reproducible large-scale platforms further limits systematic evaluation, making it difficult to test under what specific conditions such collective dynamics genuinely arise.
@@ 9. Conclusion
Acknowledgements
This research is partially sponsored by the NSF CISE grants 2302720 and 2312758, an IBM faculty award, and a CISCO research grant in Edge AI.

### v4

@@ 6. Multi-LLMGA Framework
In this section, we extend the single agent framework to multi-agent settings. Designing a multi-agent system in games is different from generic multi-agent systems because games impose unique constraints: states are partially observable, communication channels are often bandwidth-limited, and in certain scenarios direct memory sharing is disallowed to preserve realistic simulation (zhang2023building, ). To analyze how existing work addresses these challenges, we distinguish two complementary dimensions.
At the agent level, we examine how agents exchange information and integrate it into their decision-making. Communication protocols specify what messages are generated (e.g., observations, beliefs, or intentions) and how they are interpreted by receivers. At the organization level, we study three aspects: the topology of connections that shape communication flow, the allocation of tasks and roles that governs functional division of labor, and the mechanisms for ensuring scalability and robustness as groups expand. Figure 8 presents the structure of this section of different components within the multi-LLMGA system.
Figure 8. Overview of the multi-LLMGA framework.
@@ 6.2. Organizational Structure
Organizational structure defines how agents are arranged and coordinated within a multi-agent system, including the topology of their connections, the allocation of tasks and roles, and the mechanisms that ensure scalability and stability as the population grows.
@@ 8.4. Multi-LLMGA System
Prior studies have demonstrated the potential of multi-agent systems in large-scale simulations, where agents exhibit emergent behaviors. However, current large-scale multi-agent simulations remain constrained by structural and methodological limitations. Many “emergent” phenomena, such as role differentiation, norm formation, or collective planning, are closely tied to task initialization and rule design. In practice, agents are often seeded with shared goals, cooperation-oriented prompts, or predefined role templates that guide subsequent division of labor and coordination patterns. Prior studies of multi-agent societies have shown that such structural priors are widespread, from small-scale social environments (park2023generative, ) to hierarchical and large-scale simulations (sagents, ; zhao2024hierarchical, ; sid, ), where coordination often reflects the constraints of task setup rather than fully autonomous self-organization. Moreover, the lack of open and reproducible large-scale platforms further limits systematic evaluation, making it difficult to test under what specific conditions such collective dynamics genuinely arise.